mr. green is a teacher that has a job at a school , he earns k560 per hour at this school and earns a k70 at his part time job per hour. the previous week mr eaned a k3500 in both after working for 20 hours. write down the simultanaous equation that represent the above information and how long did mr greeen work on his full time job and part time job respectively


1
Expert's answer
2021-11-17T10:03:54-0500

The simultaneous equations are as follow:

"560x+70y=3500 ----(i)"

"x+y=20 ----(ii)"

subs (ii) in (i)

"560x+70(20-x)=3500"

"560x+1400-70x=3500"

"490x=2100"

"x=\t\n \\frac{2100}{490}=4.3"

"y=20-4.3=15.7"


This means Mr. Greeen did work approximately 4.3 hours on his full time job and 15.7 hours on part time job respectively during the previous week.